OpenStudy (anonymous):

Convert to proper fraction or mixed number (do not reduce to lowest terms) .8

OpenStudy (anonymous):

8/10

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

\[k = (k _{1},k _{2},k _{3})\]

OpenStudy (anonymous):

\[k_{1} = U_{x}\] \[k_{2} = U_{y}\] \[k_{3} = U_{z}\] derive x value of the middle line in the rotation matrix using: \[V[0], k[0], (k ^{\tau} V), ([k]\times V)[0]\] \[(0)\cos \theta + (-U_z)\sin \theta + (U_x)(U_y)(1- \cos \theta)\] \[= U_{x}U_{y}(1-\cos \theta) - U_{z}\sin \theta\] derive y value of the middle line in the rotation matrix using: \[V[1], k[1], (k ^{\tau} V), ([k]\times V)[1]\] \[(1)\cos \theta + (0)\sin \theta + (U_y)(U_y)(1- \cos \theta)\] \[= \cos \theta + U_{y}^2(1-\cos \theta)\] derive z value of the middle line in the rotation matrix using: \[V[2], k[2], (k ^{\tau} V), ([k]\times V)[2]\] \[(0)\cos \theta + (U_x)\sin \theta + (U_z)(U_y)(1- \cos \theta)\] \[= U_{z}U_{y}(1-\cos \theta) + U_{x}\sin \theta\]